OpenStudy (anonymous):

Write the first four terms of the sequence: a1 = 2, a2 = 3 an= an-1 * an-2

OpenStudy (anonymous):

please restate the question.

OpenStudy (anonymous):

it's written like that

jimthompson5910 (jim_thompson5910):

an= an-1 * an-2 translates to "next term = (previous term)*(term before previous term)"

jimthompson5910 (jim_thompson5910):

so the next term a3 is a3 = a2*a1 a3 = 3*2 a3 = 6 ------------------- the next term a4 is a4 = a3*a2 a4 = 6*3 a4 = 18 ------------------ and so on....